Aki's Bicycle Design has determined that when x hundred bicyles are built, the average cost per bicycle is given by C(x)=0.2x^2-0.5x+8.256, where C(x) is in hundreds of dollars. How many bicycles should the shop build to minimize the average cost per bicycle?
Answer by stanbon(75887)

Aki's Bicycle Design has determined that when x hundred bicyles are built, the average cost per bicycle is given by C(x)=0.2x^2-0.5x+8.256, where C(x) is in hundreds of dollars. How many bicycles should the shop build to minimize the average cost per bicycle?
------------
a = 0.2 ; b = -0.5 ; c = 8.256
------------
minimum cost occurs when x = -b/2a = 0.5/(2*0.2) = 0.5/0.4 = 1.25
---
Ans: 125 bikes
==============
Cheers,
Stan H.
================